let the digits of a three digit number be x,y and z. Then the value of the number is 100x+ 10y + x. Show that the difference between the number and reversed number is 99
Answer by stanbon(75887)

let the digits of a three digit number be x,y and z. Then the value of the number is 100x+ 10y + z. Show that the difference between the number and reversed number is 99
-------
reversed number: 100z + 10y + x
-----------------------------------
Equation:
(100x+ 10y + z) - (100z + 10y + x) = 99x - 99z = (99)(x-z)
============
Example:
586 - 685 = -101
200-002 = 198
-----
Note: The claim made is not true as stated.
----------------
Cheers,
Stan H.